Question 3(a)

It is known that applicants for a new graduate MBA Programme in Oxford School of Graduate Studies either get admitted or not. In the Oxford School of Graduate Studies, an admission rate of 75% has been reported. If 10 applicants are selected at random, find the probability that:

i. More than 7 applicants will be admitted.

ii. At most 7 applicants will be admitted.
Answer #1

n = 10

P = 0.75

It is a binomial distribution.

P(X = x) = nCx * px * (1 - p)n - x

i) P(X > 7) = P(X = 8) + P(X = 9) + P(X = 10)   

= 10C8 * (0.75)^8 * (0.25)^2 + 10C9 * (0.75)^9 * (0.25)^1 + 10C10 * (0.75)^10 * (0.25)^0 = 0.5256

ii) P(X < 7) = 1 - P(X > 7)

= 1 - 0.5256

= 0.4744